THE FANTASIO SET-UP (I): DESCRIPTION AND EXTENSION TOWARDS FEMTO-FANTASIO

K. DIDRICHE, C. LAUZIN, X. DE GHELLINCK, P. MACKO, A. RIZOPOULOS, P. VAN POUCKE, \underline{M. HERMAN}, Service de Chimie quantique et Photophysique CP160/09, Faculté des Sciences, Université Libre de Bruxelles (U.L.B.), Av. Roosevelt, 50, B-1050, Bruxelles, Belgium; S. KASSI, Laboratoire de Spectrométrie Physique, Laboratoire de Spectromtrie Physique, Université Joseph Fourier de Grenoble, B.P. 87, 38402 Saint-Martin-d’Heres Cedex, France.

We have built the FANTASIO set-up (for ”Fourier trANsform, Tunable diode and quadrupole mAss spectrometers interfaced to a Supersonic expansIOn”) to investigate jet-cooled molecules and dimers^{a}. The set-up will be described. New developments are on the way, to be illustrated by preliminary results. These include the doubling of the vacuum pumping system efficiency. They will also allow for a Femto OPO system to be used as a broadband, cavity enhanced absorption source to be interfaced to the supersonic expansion and to the high resolution FTIR instrument of FANTASIO.

aM. Herman, K. Didriche, D. Hurtmans, B. Kizil, P. Macko, A. Rizopoulos and P. Van Poucke, Mol. Phys., \underline{105}, 843 (2007).